# AI-Based Crypto Tokens: The Illusion of Decentralized AI?

### v1

@@ 1 Introduction
2 The Convergence of Blockchain and AI
@@
In the following section, we introduce the major AI-based tokens that currently operate within the broader cryptocurrency ecosystem, referring to them collectively as part of the decentralized artificial intelligence (DEAI) landscape. These include RENDER, a decentralized network for GPU computing [4]; Bittensor, a peer-to-peer platform for AI model collaboration [10]; Fetch.ai, which enables autonomous AI agents for real-world coordination tasks [7]; SingularityNET, a decentralized marketplace for AI services [5]; and Ocean Protocol, which facilitates decentralized data exchange for AI and analytics applications [6], among others. Figure 1 presents the market capitalization of leading existing AI-based tokens in comparison to Ethereum. The figure shows that the top four AI-based tokens exhibit relatively similar market capitalizations to each other; however, their valuations remain significantly lower than that of Ethereum.
@@ 3 Review of Major AI-Token Projects
This section reviews several prominent projects that illustrate the AI-token concept in practice. Each project is analyzed based on its platform architecture, the function and utility of its token, the consensus or trust mechanism it employs, and its business or application model. Table 1 presents a structured comparison of these projects, highlighting key aspects such as their underlying blockchain networks, their roles, token utilities, consensus mechanisms, and business models.
@@ 3.6 Merging FET, AGIX, and OCEAN
Beyond those major AI-token platforms, several other projects contribute to the broader ecosystem by targeting specialized niches at the intersection of AI and blockchain. Numeraire (NMR) [8] incentivizes accurate financial predictions through a stake-based competition model, turning algorithmic performance into a tradable asset. Story (IP) [54] introduces a peer-to-peer blockchain protocol for registering, exchanging, and monetizing intellectual property as programmable assets, enabling decentralized, automated licensing and collaboration across AI agents, software, and users through a multi-core architecture and cryptoeconomic incentives. Cortex (CTXC) [47] pioneered on-chain AI inference by embedding machine learning capabilities within smart contracts, though it has seen limited adoption. Artificial Liquid Intelligence (ALI) [48] powers Alethea AI’s intelligent NFTs, combining generative AI with token-based governance to manage ethical content moderation. DeepBrain Chain (DBC) [49], one of the earliest attempts at decentralized AI compute, aimed to create a distributed GPU marketplace but struggled to scale against newer competitors. Their diverse architectures and use cases provide valuable insights into how token economies can be structured to support different facets of decentralized AI, while also highlighting recurring challenges such as adoption barriers, scalability, ethical concerns, and the need for sustainable incentive mechanisms.
Note that this paper focuses solely on tokens that explicitly identify themselves as AI tokens. However, there are several notable projects that, while primarily addressing core blockchain challenges, have recently shifted their focus to AI. For instance, P2P storage solutions like IPFS, Arweave, and ICP [50, 51], which serve as the backbone of decentralized storage, are now also working to accommodate decentralized AI. Similarly, The Graph [52, 53], often referred to as the "Google of blockchain" due to its decentralized indexing capabilities, has expanded its focus to include AI by publishing their new whitepaper "The Graph as AI Infrastructure". Chainlink [74], a prominent oracle solution in the blockchain ecosystem, is now developing AI-driven oracle agents. Even some Layer 1 blockchain networks, such as Near, have begun shifting their attention to AI-focused initiatives.To conclude this section, Table 2 presents the classification of existing AI-based tokens within the cryptocurrency ecosystem.
@@ 4.4 Ecosystem Bootstrapping and Network Effects
4.5 Governance and Evolution Challenges
@@ 4.6 Delivering Value Beyond Centralized AI
Given the challenges outlined above, it is understandable why current AI-token ecosystems have struggled to deliver value that clearly surpasses centralized AI alternatives. At present, many decentralized solutions involve greater complexity, slower performance, and limited usability. As a result, the theoretical benefits of decentralization such as censorship resistance, transparent governance, and collective ownership have not yet proven compelling enough to drive mass adoption outside of ideologically motivated communities or early adopters.
@@ 5 Future Directions and Opportunities
Despite current limitations, the intersection of AI and blockchain continues to evolve, with several promising developments on the horizon. Advances in on-chain verifiability (e.g., zkML [72]), improved incentive systems [96], privacy-preserving computation, and cross-chain interoperability are paving the way for more trustless and scalable AI networks. New paradigms such as decentralized collaborative model training, compute-to-data frameworks, and modular AI ecosystems may enable AI-token projects to deliver unique value beyond what centralized systems can offer. These directions, supported by active research and experimentation, represent meaningful steps toward more open, intelligent, and community-driven AI infrastructures.
@@ 5.1 Verifiable Off-Chain Computation and AI Oracles
A major step toward trustless decentralized AI is enabling on-chain verification of off-chain computation. Zero-knowledge proofs for machine learning (ZKML [72]) allow an AI provider to submit not only an output but also a cryptographic proof that the output was generated by a specific model on a given input. Although currently limited to lightweight models, ongoing research and hardware acceleration may soon make this practical for broader AI use. Trusted Execution Environments (TEEs) [75] provide an alternative, enabling secure model execution within hardware enclaves and generating attestations that can be verified on-chain. This shifts trust from server operators to hardware providers and is already being explored in federated learning contexts. Additionally, AI oracles, such as those under development by Chainlink [73, 74], propose a model in which multiple nodes independently execute an AI task and report only when a consensus is reached. This quorum-based approach could make AI inference verifiable and composable within smart contracts, paving the way for trust-minimized AI services accessible on-chain.
@@ 5.2 Specialized Blockchains and Layer-2s for AI
Emerging infrastructure tailored for AI workloads may address current performance bottlenecks in decentralized networks. AI-dedicated Layer-1s could incorporate native GPU support, larger block sizes, and sharding by AI task type (e.g., language or vision), enabling scalable, parallelized inference and training. Novel consensus mechanisms like Proof-of-Useful-Work [77, 76], where miners perform machine learning computations, could further align network security with productive AI contributions, though this depends on solving verification challenges. Alternatively, Layer-2 solutions [78, 79, 38, 80, 81], for instance, Golem (GLM) [83] iExec (RLC) [84], which focuses on decentralized GPU computing, has migrated from Ethereum Layer 1 to Ethereum’s Layer 2. In partucular, zk-rollups, offer a promising path. By batching off-chain AI model executions and submitting zero-knowledge proofs to Layer-1s, rollups can significantly reduce costs and enable verifiable inference at scale. Projects like ZkMatrix [82] hint at this direction.

### v2

@@ 1 Introduction
2 Related Work
In this section, we review existing literature on the convergence of artificial intelligence and blockchain technologies, followed by a discussion on the illusion of decentralized AI and the speculative nature of crypto token economies.
2.1 The Convergence of Blockchain and AI
@@
2.2 Decentralization Illusions and Speculative Token Economies
The term “illusion of decentralized AI” refers to cases in which blockchain-based AI projects present a decentralized architecture in theory, while retaining centralized control over core operations in practice. This perception is often maintained through structural features commonly associated with decentralization such as token-based governance, geographically distributed nodes, or autonomous project branding. Despite the fact that critical components like model training, data hosting, and protocol updates remain under the authority of a limited group of insiders. In such arrangements, decentralization exists in name only, as substantive control over decision-making, infrastructure, and economic flows remains concentrated among project founders, venture capital entities, or centralized service providers.
This decentralization illusion is not unique to AI-based tokens. Across the broader Web3 ecosystem, similar patterns have emerged where decentralization is assumed by design but fails in implementation. Researchers have increasingly emphasized the need for rigorous, quantifiable frameworks to assess whether authority, participation, and infrastructure are genuinely decentralized. In the absence of deliberate mechanisms to prevent concentration of control, blockchain-based systems risk replicating traditional power hierarchies while merely presenting an appearance of decentralization [39, 40]. In the context of decentralized AI, this means that tokenization or blockchain use alone is insufficient. Genuine decentralization must be evaluated based on who governs the models, how infrastructure is provisioned, and whether the community can meaningfully participate in development and oversight. The “illusion of decentralized AI” thus captures a critical gap between the ideals of Web3 and the realities of control in current AI-token ecosystems.
This conceptual gap is compounded by the fact that many so-called utility tokens have failed to deliver on their promised functionality, instead becoming vehicles for speculative trading. Empirical studies consistently show that the majority of utility tokens experience disproportionately high levels of speculation compared to actual platform usage [42, 41, 43]. A comprehensive analysis of 891 Ethereum-based tokens revealed a 90% decline in average utility usage since 2017, even as speculative activity surged particularly following the rise of decentralized finance (DeFi), which introduced leveraged trading mechanisms that further shifted attention away from functional use [41]. This trend spans multiple sectors, from decentralized storage to gaming, where token holders increasingly prefer to trade or hoard tokens in anticipation of price appreciation, rather than use them for access or services as originally intended.
This disconnect is not merely behavioral but often embedded in token design. By making access tokens freely tradable, many projects inadvertently prioritize financial incentives over organic platform engagement. The Internet Computer Protocol (ICP) serves as a vivid example: despite initial hype and a multi-billion dollar valuation, its token collapsed by over 95% shortly after launch, with real user adoption stagnating [43]. Similar dynamics have been observed in Filecoin, where token value has been driven more by mining and speculation than by genuine demand for decentralized storage [44]. These cases underscore a broader trend in tokenized ecosystems: despite ambitious narratives, many tokens have not fulfilled their intended utility roles, as speculative behavior systematically overrides platform use.
@@ 3 Review of Major AI-Token Projects
This section examines several leading projects that exemplify the AI-token paradigm, with a focus on their platform architecture, token functionality, consensus mechanisms, and business or application models, as summarized in Table 1. The selected projects are drawn from the top-ranked AI tokens by market capitalization, according to CoinMarketCap [^1] https://coinmarketcap.com/view/ai-big-data/ and CoinGecko [^2] https://www.coingecko.com/en/categories/artificial-intelligence, and are limited to those that explicitly identify themselves as AI tokens. Figure 1 presents the market capitalizations of the top AI-based tokens, with Ethereum included for comparative context. As shown, the top four AI tokens exhibit comparable valuations to one another; however, their market capitalizations remain significantly lower than that of Ethereum. This selection criterion ensures the analysis remains focused on initiatives that are AI-native by both design and narrative.
It is important to note that several prominent blockchain projects, although not originally developed as AI platforms, are now categorized under the AI sector on platforms such as CoinMarketCap and CoinGecko, reflecting recent strategic shifts toward AI applications. These include peer-to-peer storage networks such as IPFS, Arweave, and the Internet Computer Protocol (ICP) [45, 46], which are increasingly used for decentralized AI data and model hosting. The Graph [47, 48], often referred to as the “Google of blockchain” for its decentralized indexing capabilities, has also rebranded its roadmap to support AI-related infrastructure, as outlined in its recent whitepaper “The Graph as AI Infrastructure.”. Chainlink [85], a leading oracle network, is actively developing AI-driven oracle agents, and Layer 1 blockchain protocols such as Near [49] have launched dedicated initiatives aimed at supporting AI-native development. While these projects reflect the growing convergence between blockchain and AI, they are excluded from the core analysis in this paper, as they do not represent tokens that were initially or explicitly positioned as AI-focused. Their inclusion in the AI category, however, signals the expanding scope and appeal of decentralized AI as a frontier across the broader Web3 ecosystem.
In this section, we review the top AI-based tokens include RENDER, a decentralized network for GPU computing [4]; Bittensor, a peer-to-peer platform for AI model collaboration [10]; Fetch.ai, which enables autonomous AI agents for real-world coordination tasks [7]; SingularityNET, a decentralized marketplace for AI services [5]; and Ocean Protocol, which facilitates decentralized data exchange for AI and analytics applications [6], among others.
@@ 3.6 Merging FET, AGIX, and OCEAN
Beyond those major AI-token platforms, several other projects contribute to the broader ecosystem by targeting specialized niches at the intersection of AI and blockchain. Numeraire (NMR) [8] incentivizes accurate financial predictions through a stake-based competition model, turning algorithmic performance into a tradable asset. Story (IP) [61] introduces a peer-to-peer blockchain protocol for registering, exchanging, and monetizing intellectual property as programmable assets, enabling decentralized, automated licensing and collaboration across AI agents, software, and users through a multi-core architecture and cryptoeconomic incentives. Cortex (CTXC) [58] pioneered on-chain AI inference by embedding machine learning capabilities within smart contracts, though it has seen limited adoption. Artificial Liquid Intelligence (ALI) [59] powers Alethea AI’s intelligent NFTs, combining generative AI with token-based governance to manage ethical content moderation. DeepBrain Chain (DBC) [60], one of the earliest attempts at decentralized AI compute, aimed to create a distributed GPU marketplace but struggled to scale against newer competitors. Their diverse architectures and use cases provide valuable insights into how token economies can be structured to support different facets of decentralized AI, while also highlighting recurring challenges such as adoption barriers, scalability, ethical concerns, and the need for sustainable incentive mechanisms. To conclude this section, Table 2 presents the classification of existing AI-based tokens within the cryptocurrency ecosystem.
@@ 4.4 Ecosystem Bootstrapping and Network Effects
4.5 Governance and Project Evolution Challenges
@@
4.6 Accountability and Legal Responsibility in Decentralized AI
While decentralized AI-token projects promise openness, autonomy, and trustless execution, they also raise pressing concerns around accountability and legal responsibility. Traditional AI systems typically operate under a central provider who can be held liable for malfunctions or harm. In contrast, decentralized platforms like Fetch.ai, Ocean Protocol, and SingularityNET distribute control across pseudonymous actors, making it difficult to determine who, if anyone, bears responsibility in the event of failure or misuse. As recent legal scholarship points out, these systems often create a “liability vacuum,” where no single party is clearly accountable under existing legal frameworks  [^3] https://www.blockchainandthelaw.com/2023/04/dao-deemed-general-partnership-in-negligence-suit-over-crypto-hack-prompting-decentralized-orgs-to-rethink-corporate-formation/.
Several high-profile projects attempt to mitigate this gap through hybrid governance models or foundation structures. For instance, Ocean Protocol and Fetch.ai maintain formal legal entities that can serve as regulatory touchpoints, while projects like Render and Bittensor rely on DAO-based voting mechanisms for internal oversight. However, token-based governance does not imply legal personhood, and scholars have raised concerns that DAO participants or token holders could, under certain conditions, be deemed members of unincorporated partnerships, thereby exposing them to liability. This has been exemplified by the U.S. court ruling in the Ooki DAO case [^4] https://www.cftc.gov/PressRoom/PressReleases/8715-23, which signaled that decentralized participation may not exempt individuals from legal accountability.
Technical solutions such as proof-of-quality staking mechanisms (e.g., Numeraire, Bittensor) or auditability through blockchain-based provenance have been proposed to enhance internal accountability. Yet these remain voluntary and do not fulfill the legal requirements of redress, especially under emerging global AI regulations. The EU AI Act [69], for example, imposes obligations such as transparency, data governance, and post-deployment monitoring responsibilities that decentralized AI networks may struggle to assign without a designated operator.
As legal and regulatory frameworks evolve, AI-token ecosystems must address the mismatch between distributed operation and centralized legal responsibility. Without clearer outline of liability whether through DAO registration, delegated entities, or protocol-level insurance, projects risk undermining user trust and exposing contributors to unforeseen legal risks. Accountability in decentralized AI will require more than on-chain governance; it demands structural innovations that bridge blockchain design with real-world legal norms.
@@ 4.7 Delivering Value Beyond Centralized AI
While this study focuses on AI-based tokens, similar challenges have emerged across a broader range of blockchain projects, suggesting that the underperformance of utility tokens is often a systemic issue. During the 2017–2018 ICO boom, numerous projects across sectors such as IoT, cloud storage, and social media launched tokens with ambitious promises but failed to achieve meaningful user adoption. Empirical studies [41, 42, 43] show that token valuations frequently decouple from actual usage, indicating that speculative trading often outweighs real utility. Helium, for example, attracted substantial investment and deployed widespread infrastructure, yet by 2022, its IoT data network generated only minimal revenue, highlighting a disconnect between token-driven supply growth and end-user demand  [^5] https://cointelegraph.com/news/critique-on-helium-s-6-5k-monthly-revenue-causes-a-stir.
A similar pattern is evident in Filecoin, which built massive decentralized storage capacity but saw minimal utilization in its early years, with just 3.8% of space filled compared to 40–70% for centralized providers [44]. Despite incentives and subsequent efforts to boost demand, adoption remained slow. These cases reveal a common issue in tokenized ecosystems: while token rewards can effectively stimulate supply, generating sustained demand for decentralized services, particularly when competing with established centralized alternatives, proves far more difficult. Such outcomes underscore that the limitations observed in AI-token projects are not unique, but reflective of broader structural challenges in aligning token models with real-world utility. These examples from outside the AI domain illustrate that token underperformance is frequently a matter of lack of product-market fit and network effects, rather than something intrinsic to AI.
Given the challenges outlined in Sec. 4, it is understandable why current AI-token ecosystems have struggled to deliver value that clearly surpasses centralized AI alternatives. At present, many decentralized solutions involve greater complexity, slower performance, and limited usability. As a result, the theoretical benefits of decentralization such as censorship resistance, transparent governance, and collective ownership have not yet proven compelling enough to drive mass adoption outside of ideologically motivated communities or early adopters.
@@
Table 4: Summary of Key Achievements and Shortcomings of Selected AI Token Projects
Project Token Key Achievements Notable Shortcomings
FET (Fetch.ai) Launched a Cosmos-based mainnet enabling agent-based coordination and WASM smart contracts. Most AI tasks run off-chain. Governance is still significantly influenced by the foundation, raising centralization concerns.
RNDR (Render) Enabled a decentralized GPU rendering network with measurable growth (e.g., 121% QoQ usage increase in Q2 2023)a. Adopted a burn-and-mint model and DAO-based community governance. Most rendering tasks are conducted off-chain, requiring trust in node integrity. Initial development was centralized via OTOYb, and current focus remains narrow (e.g., graphics rendering).
TAO (Bittensor) Pioneered mechanism where model quality determines block rewards [10]. Established a peer-evaluated AI model marketplace and introduced subnet specialization. Scalability remains a challenge, as verifying model quality requires peer coordination. The outputs are not state-of-the-art, and participation barriers (hardware, ML expertise) limit decentralization.
OCEAN (Ocean Protocol) Introduced Compute-to-Datac, allowing algorithms to run on private datasets without revealing raw data. Enabled data tokenization for monetization and supported DeSci integrations. The marketplace sees limited real demand. Many datasets are test entries or unused, and competition from free, centralized alternatives (e.g., Kaggle) undercuts its utility proposition.
AGIX (SingularityNET) Built a decentralized AI service marketplace and led the ASI Alliance merger (AGIX, FET, OCEAN). Continued to fund AGI-related research (e.g., OpenCog Hyperon). Platform usage is low, with many listed services seeing no traffic. Trust mechanisms for service quality are lacking, and multi-chain operations introduce technical complexity.
CTXC (Cortex) Implemented on-chain model inference via the Cortex Virtual Machine (CVM) with GPU-backed PoW consensus. Despite the novel architecture, the project has seen minimal developer engagement, and scaling inference fully on-chain remains infeasible with current resources.
NMR (Numeraire) Enabled crowdsourced model predictions for a hedge fund, using stake-weighted incentives to reward accuracy. The platform is heavily centralized in its model selection process, and broader adoption beyond the Numerai ecosystem is limited.
IP (Story) Proposed a blockchain-based licensing system for AI-generated and co-created IP, enabling programmable royalties. The ecosystem remains underdeveloped, with little visibility or user adoption, and limited clarity on execution scalability.
DBC (DeepBrain Chain) Offered a decentralized GPU marketplace for AI compute, utilizing a Substrate chain and validator staking. The project has suffered from poor documentation and declining community activity. Verifiability of output remains unsolved.
ALI (Alethea AI) Pioneered interactive NFTs powered by AI personas and showcased avatar monetization via AI dialogue systems. The use case is narrow and largely speculative, with limited mainstream demand or technical innovation in AI processing.
  • a
    https://www.coinfeeds.ai/crypto-blog/render-network
  • b
    https://home.otoy.com/the-company/
  • c
    https://docs.oceanprotocol.com/developers/compute-to-data
@@ 5 Future Directions and Opportunities
Despite current limitations, the intersection of AI and blockchain continues to evolve, with several promising developments on the horizon. Advances in on-chain verifiability (e.g., zero-knowledge proofs such as zkML [80], zkPoT [81], [82, 83]), improved incentive systems [108], privacy-preserving computation, cross-chain interoperability, composability, and modular blockchain architectures are paving the way for more trustless and scalable AI networks. Supported by active research and experimentation, these developments represent meaningful steps toward building open, intelligent, and community-driven AI infrastructures that can deliver value beyond what centralized systems are capable of.
@@ 5.1 Verifiable Off-Chain Computation and AI Oracles
A major step toward trustless decentralized AI is enabling on-chain verification of off-chain computation. Zero-knowledge proofs for machine learning (zkML [80]), for deep learning (zkPoT [81]) allow an AI provider to submit not only an output but also a cryptographic proof that the output was generated by a specific model on a given input. Although currently limited to lightweight models, ongoing research and hardware acceleration may soon make this practical for broader AI use. Trusted Execution Environments (TEEs) [86] provide an alternative, enabling secure model execution within hardware enclaves and generating attestations that can be verified on-chain. This shifts trust from server operators to hardware providers and is already being explored in federated learning contexts. Additionally, AI oracles, such as those under development by Chainlink [84, 85], propose a model in which multiple nodes independently execute an AI task and report only when a consensus is reached. This quorum-based approach could make AI inference verifiable and composable within smart contracts, paving the way for trust-minimized AI services accessible on-chain.
@@ 5.2 Specialized Blockchains and Layer-2s for AI
Emerging infrastructure tailored for AI workloads may address current performance bottlenecks in decentralized networks. AI-dedicated Layer-1s could incorporate native GPU support, larger block sizes, and sharding by AI task type (e.g., language or vision), enabling scalable, parallelized inference and training. Novel consensus mechanisms like Proof-of-Useful-Work (PoUW) [88, 87], where miners perform machine learning computations, could further align network security with productive AI contributions, though this depends on solving verification challenges. Academic proposals like Coin.AI [89] have outlined mechanisms where block validity depends on a trained model meeting performance thresholds, with verification handled efficiently by the network. A real-world example is Bittensor which rewards miners based on the value their models contribute to a shared AI ecosystem. If implemented securely, PoUW could align blockchain incentives with the generation of valuable AI artifacts, transforming blockchain networks into decentralized compute resources.
Alternatively, Layer-2 solutions [90, 91, 38, 92, 93], for instance, Golem (GLM) [95] iExec (RLC) [96], which focuses on decentralized GPU computing, has migrated from Ethereum Layer 1 to Ethereum’s Layer 2. In partucular, zk-rollups, offer a promising path. By batching off-chain AI model executions and submitting zero-knowledge proofs to Layer-1s, rollups can significantly reduce costs and enable verifiable inference at scale. Projects like ZkMatrix [94] hint at this direction.
5.3 Modular Blockchain Architectures for Decentralized AI
A key architectural shift shaping the future of decentralized AI is the transition from monolithic to modular blockchain designs. Traditional monolithic blockchains (e.g., Bitcoin, Ethereum (pre-rollups)) require each node to perform all core functions consensus, data availability, and transaction execution which limits scalability and makes it difficult to accommodate computation heavy tasks such as AI model training or inference. In this design, every node must process every transaction, creating performance bottlenecks that hinder the integration of advanced AI workloads. Modular blockchains (e.g., Cosmos, Celestia, Ethereum (with rollups)), by contrast, decouple these functions into separate layers or components. For example, one layer might focus solely on transaction ordering and consensus, while a separate execution layer handles the state changes and computational logic. This division of labor allows each component to be optimized independently, improving efficiency and throughput while preserving system wide coherence.
For decentralized AI, the implications of modularity are particularly significant. AI applications often require flexible, high-throughput environments to support data processing, model execution, and real-time responsiveness. Modular architectures enable the deployment of specialized execution layers tailored to AI-specific tasks, without placing additional strain on the base chain. This model not only enhances scalability but also encourages composability where different components of the AI stack (e.g., data access, computation, governance) can operate on separate yet interoperable chains. Projects such as the Artificial Superintelligence Alliance, which includes Fetch.ai (FET) and SingularityNET (AGIX), illustrate this direction by envisioning interoperable AI modules that can operate across a network of purpose built chains. Rather than waiting for a single blockchain to meet all computational and coordination needs, modularity allows AI developers to build scalable systems by composing specialized layers, each responsible for a distinct function. This paradigm shift represents a more adaptable and sustainable path forward for decentralized AI infrastructure.
@@ 5.8 Ethical and Equitable AI Ecosystems
5.9 Guidelines for Launching a Decentralized AI Token
Drawing on the patterns observed in our review summarized in Table LABEL:tab:summary_achievement_shortcomings, which highlights the key strengths and persistent limitations of leading AI-token projects, it becomes evident that many initiatives face common design and adoption challenges. These findings underscore the importance of critically assessing the rationale and structure behind launching a new token especially AI-based token. To support more sustainable and purpose-driven deployments, the following framework offers practical guidelines for evaluating whether, and under what conditions, an AI-based token should be introduced.
  • Validate the Need for Decentralization: Assess whether decentralization is essential to the project’s goals [120]. If the AI service can operate effectively under a centralized model, the introduction of a token may introduce unnecessary complexity. Tokens should only be deployed where decentralization provides clear benefits, such as enhanced data ownership, censorship resistance, or trustless coordination.
  • Establish Product Market Fit: Before launching a token, ensure that the core AI product addresses a well-defined user need or pain point. Product market fit should be validated through user feedback, prototype testing, or early adoption metrics. Many decentralized AI projects risk failure by attempting to replicate centralized models without sufficient evidence that decentralization improves the user experience or solves a meaningful problem. Tokens should not be introduced merely as a funding mechanism or technical novelty; instead, their design and function must be grounded in a service that delivers clear and differentiated value to its target users. Without product–market fit, token demand will be speculative at best and unsustainable in the long term.
  • Define Clear Token Utility: Specify the token’s function within the platform. Tokens should serve a critical role such as enabling access to AI services, facilitating payments, or supporting governance. A token that lacks unique, indispensable utility risks redundancy and diminished value. Committing to accept only the native token for core services can help anchor its economic relevance.
  • Align Incentives and Promote Network Effects: Design token mechanics that incentivize productive contributions (e.g., data sharing, model development, compute provisioning). The token should help coordinate stakeholder behavior and amplify network growth. However, care must be taken to avoid speculative incentives that detract from the platform’s actual functionality [121].
  • Assess Technical Readiness: Ensure that the AI infrastructure is sufficiently developed to support meaningful token use. Launching a token prematurely before services are usable or integrated with the token can damage credibility. Consider technical factors such as workload distribution (on-chain vs. off-chain) and the feasibility of token integration with core workflows.
  • Consider Regulatory and Governance Implications: Evaluate the legal and governance frameworks surrounding the token. Determine whether the token might be classified as a security or financial instrument in key jurisdictions, and structure the launch to meet decentralization thresholds where applicable. Additionally, prepare for community led governance, including treasury management and protocol upgrades, before transferring control of the ecosystem to token holders.
This structured approach helps ensure that token launches within decentralized AI systems are grounded in genuine utility, supported by technical maturity, and guided by responsible governance.